# No Bells, Just Whistles: Sports Field Registration by Leveraging Geometric Properties

### v2

@@ 4 Experiments
We implement our proposed method for sports field registration. This section provides an overview of the datasets utilized, the evaluation metrics employed to assess our approach, and implementation specifics. Subsequently, we present both qualitative and quantitative results, comparing our method with state-of-the-art approaches.
4.1 Datasets
@@
To evaluate our method, we follow state-of-the-art methods by using the SoccerNet Calibration [8], the WorldCup [21] and the TS-WorldCup [7] soccer datasets.
@@
SN-Calib Dataset: The SoccerNetV3-Calibration (SN23) dataset [8] comprises 22,816 images extracted from SoccerNet [12] videos and encompasses a broadcast-based multi-view nature, offering a broader range of camera perspectives beyond the main broadcast camera. Cioppa et al. [8] provide annotations for all segments of the soccer field, encompassing lines, conics and goal posts. For each visible segment on the court, at least two annotated positions are provided, optimally representing the segment in a polyline format. For the conics drawn on the pitch, the annotations consist of a list of points that roughly give the circle shape when connected. Additionally, Theiner et al. [33] provided manual camera view annotations for the SoccerNetv3-Calibrationim 2022 (SN22) dataset version, allowing the creation of data subsets taking into account the camera view distribution.
@@ 4.2 Evaluation Metrics
The quality of estimated camera parameters or homography matrices can be evaluated both in 2D image and world space.
[…]
@@
Intersection over Union (IoU): The intersection over union (IoU) includes two components: $\text{IoU}_{part}$ quantifies the visible area of the video frame by warping the video frame using both the refined homography and the GT homography, projecting them onto the template, and then calculating the IoU. $\text{IoU}_{whole}$ evaluates the entire sports field by warping the template with the refined homography, projecting it onto the original template, and calculating the IoU.
@@
Projection Error: The projection error was quantified as the average distance, in meters, between the projected points using the predicted homography and the GT homography. To achieve this, we uniformly sampled $2,500$ pixels from the visible field area of the camera image and projected them onto the field to compute the distance. The standard dimensions of a soccer field are $105\times 68$ meters.
@@
Reprojection Error: The reprojection error was calculated by averaging the distance between the reprojected points in the video frame, utilizing both the predicted and the GT homography.
4.3 Implementation Details
@@
Due to the absence of publicly available results on the multiple-view SN23 distribution, we trained two models from scratch: Multi-view (MV) and Single-view (SV). The latter is composed almost entirely of non-replay frames, ensuring a high percentage of central camera shots. We train separate networks for the keypoints and line extremities detection tasks on the SN23-train dataset [8]. For the MV model, we train for $200$ epochs, using an initial learning rate of $1e^{-2}$ and a batch size of $2$. For the SV model, we train for $200$ epochs, using an initial learning rate of $1e^{-5}$ and a batch size of $2$. We utilize the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. ${l}_{2}$-loss is used for heatmap regression in both neural networks. Data augmentation such as random horizontal flip, color jitter, and Gaussian noise are applied to enhance model robustness and generalization. Furthermore, we fine-tune both SV networks on the WC14 and TSWC datasets. GT homographies are transformed into our proposed keypoint sets and line extremities by projecting their respective world coordinates to the field’s ground plane. The experiments are conducted on a single NVIDIA GeForce RTX 2080 Ti GPU with $12$ GB of memory.
4.4 Results and comparisons
@@
In this section, we present the results of an extensive evaluation, divided into camera calibration and homography estimation. The former assesses the accuracy of individual camera parameters using Acc@t metric, while the latter evaluates the quality of homography estimation using IoU metrics, projection error, and reprojection error.
4.4.1 Camera Calibration
@@
In team sports such as soccer, the action takes place on a nearly planar field. Consequently, most methods utilize homography estimation to map all elements positioned on this plane but cannot project non-planar points such as points belonging to goal posts or crossbars. Conversely, Theiner et al. [33] employs a 3D model of the soccer field to extract camera pose and intrinsic parameters directly. In homography-based approaches, parameter retrieval is accomplished through homography decomposition (HDecomp). We conduct a quantitative comparison of our proposed method with respect to state-of-the-art approaches [6, 23, 33] on the SN22-test-center dataset, comprising only images where the main camera center is visible (1,454 images). Furthermore, utilizing the SoccerNet annotation format for the WC14-test dataset provided by Theiner et al. [33], we conduct a comparison of our proposed method’s performance in camera parameter estimation on the WorldCup 2014 dataset distribution.
We report the statistics from [33] paper for the results of state-of-the-art approaches [6, 23, 33]. As shown in Tables 1-2 for SN22-test and WC14 datasets, respectively, our SV method outperforms state-of-the-art approaches on several metrics for both datasets. Minor variations in CR arise due to our approach requiring a minimum number of visible keypoints for calibration.
Finally, we evaluate our method on the entire SN23-test dataset, which defines the first public camera calibration benchmark on the SoccerNetV3-Calibration [8] dataset, and the first one extending calibration assessments to encompass multiple-view scenarios.
[…]
@@
Table 1: Evaluating camera calibration on SoccerNet distributions. We assess our Multi-view model on the entire SN23-test dataset and quantitatively compare our Single-view model with other approaches on the SN22-test-center dataset.
@@
[…]
Table 2: Quantitative comparison of our Single-view model on camera calibration conducted on the WC14-test dataset.
4.4.2 Homography Estimation
The proposed method is compared to state-of-the-art approaches [6, 23, 9, 36, 37, 27, 31, 7, 28, 33, 25] using the WC14-test dataset. Additionally, our method is also compared to state-of-the-art approaches [6, 27, 7, 25, 28] using the TSWC-test dataset. For computing IoU, projection error, and reprojection error, we adopt the approach outlined in [7]. The dimensions of the soccer field template are set at $115\times 74$ yards for fair comparison.
Regarding the evaluation on the WC14-test dataset, we present the performance metrics based on the findings reported in the respective papers of the state-of-the-art approaches, as detailed in Table 3. To ensure a fair comparison with [9], we use the results of the approach "ours-w/o-players" reported in their paper. Our fine-tuned model achieves competitive results in comparison with other methods, which are topped by [31, 28], on most of the metrics. Notably, we attain state-of-the-art performance in the median value of the reprojection error without requiring further homography refinement, unlike [6, 7, 28].
For the evaluation on the TSWC-test dataset, we report the statistics from the papers [27, 7, 25, 28] in Table 3, observing how our fine-tuned model outperforms those methods in most of the metrics, demonstrating the effectiveness of our approach without the need for further refinement.
@@ 4.4.2 Homography Estimation
Table 3: Evaluating the homography estimation on WC14-test and TSWC-test. $\ast$ denotes the methods trained on SoccerNet distribution, $\dagger$ denotes the methods fine-tuned on the WC14 dataset and $\ddagger$ denotes the methods fine-tuned on the TSWC one.
[…]
             66.8 73.9 91.9 85.9 57.4
66.9 86.0 91.9 85.9 57.5
74.6 89.1 92.1 91.8 68.6
73.8 87.8 91.1 96.5 71.3
75.3 89.4 91.1 97.8 73.7
@@
Table 4: Ablation study of our keypoint sets. The table shows the effect of every keypoint set on the SN22-test-center dataset.
4.4.3 Ablation Study on Keypoint Sets Contribution
The contribution of each keypoint set, namely $\mathcal{K}p_{e}$, $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$, and $\mathcal{K}p_{3}$, is analyzed in Table 4. The integration of $\mathcal{K}p_{1}$ notably enhances accuracy, particularly along the midfield line, as no points were available in the previous sets. The inclusion of $\mathcal{K}p_{2}$ augments the completeness rate by increasing the keypoint density within the field circles but shows a small decrease in accuracy metrics. This is due to the still low number of keypoints across these field circles; we are adding new low-quality detections, hence decreasing the overall accuracy. Furthermore, the integration of $\mathcal{K}p_{3}$ further elevates accuracy and completeness rate, ultimately defining our robust keypoints grid. These conclusions can be straightforwardly seen in the field model grid presented in Fig. 2. Although $\mathcal{K}p_{e}$ does not contribute to a score boost, its significance lies in increasing the number of points to allow the computation of the subsequent sets during the training data generation.
5 Conclusion
@@
In this paper, we introduce a novel framework for 3D sports field registration. Our proposed pipeline adopts a minimalist approach by solely utilizing the geometric properties of the soccer field. We demonstrate superior performance in 3D camera calibration on SoccerNet and WorldCup 2014 datasets compared to state-of-the-art methods, while also achieving comparable results in homography estimation on WorldCup 2014 and TS-WorldCup datasets. Additionally, we extend our pipeline to multiple-view camera calibration, thereby establishing the first public multiple-view broadcast-based camera calibration benchmark in soccer. Our method exhibits promising results, highlighting the effectiveness of utilizing a robust field template without the need for further refinements. As long as the video frame distortions are not too harsh (i.e. fisheye shots) and a minimum of four keypoints are visible, sports field registration is shown to be effective. In future work, we plan to enhance our approach by incorporating temporal consistency between subsequent video frames, aligning better with the nature of sports video broadcasts.
@@
Acknowledgment. This work has been supported by the project MoHuCo PID2020-120049RB-I00 funded by MCIN/ AEI /10.13039/501100011033.

### v3

@@ IV Experiments
This section provides an overview of the datasets we use, the evaluation metrics employed to assess our approach, as well as implementation details. Subsequently, we present both qualitative and quantitative results, comparing our method with state-of-the-art approaches and performing an ablation study of the keypoint sets effect on the method performance.
IV-A Datasets
@@
To evaluate our method, we compare our results with state-of-the-art methods on the SoccerNet-Calibration [8], the WorldCup [23] and the TS-WorldCup [7] soccer datasets.
@@
SN-Calib Dataset: The SoccerNetV3-Calibration (SN23) dataset [8] comprises 22,816 images extracted from SoccerNet [12] videos and encompasses a broadcast-based multi-view nature, offering a broader range of camera perspectives beyond the main broadcast camera. Cioppa et al. [8] provide annotations for all segments of the soccer field, encompassing lines, conics and goal posts. For each visible segment on the court, at least two annotated positions are provided, optimally representing the segment in a polyline format. For the conics drawn on the pitch, the annotations consist of a list of points that roughly give the circle shape when connected. Additionally, Theiner et al. [46] provided manual camera view annotations for the SoccerNetv3-Calibration 2022 (SN22) test set version, allowing the creation of data subsets taking into account the camera view distribution.
@@ IV-B Evaluation Metrics
The quality of estimated camera parameters or homography matrices can be evaluated both in 2D and 3D spaces.
Jaccard Index ($\text{JaC}_{\gamma}$): Following Magera et al. [28] benchmarking protocol, the evaluation relies on calculating the reprojection error between each annotated point and the line to which it belongs. Adopting a binary classification approach, each pitch segment is treated as a single entity. To be considered correctly detected, all points within the segment must have a reprojection error smaller than a threshold. The projection of pitch elements from densely sampled points of the soccer field 3D model yields a polyline for each segment. Therefore, a polyline representing a soccer field segment $s$ is classified as a true positive (TP) if $\forall p\in s:\min\left(d(p,\hat{s})\right)<\gamma$, being $\hat{s}$ the corresponding annotated segment and $\gamma$ the distance threshold in pixels. Otherwise, this segment is counted as a false positive (FP). Segments only present in the annotations are counted as false negatives (FN). Hence, the Jaccard Index for camera calibration, $\text{JaC}_{\gamma}$, at a threshold of $\gamma$ pixels is defined as:
$\text{JaC}_{\gamma}=\frac{\text{TP}_{\gamma}}{\text{TP}_{\gamma}+\text{FN}+\text{FP}},$ (8)
where it serves as a measure of calibration accuracy. We also measure the completeness rate (CR) as the number of camera parameters provided divided by the number of images with more than four semantic line annotations in the dataset. The final score (FS) as an evaluation criterion is calculated as the product of CR and $\text{JaC}_{5}$.
@@
Intersection over Union: The intersection over union (IoU) includes two components. First, $\text{IoU}_{part}$ quantifies the visible area of the video frame by warping that frame using both the estimated homography and the GT one, projecting them onto the template, and then calculating the IoU. Second, $\text{IoU}_{whole}$ evaluates the entire sports field by warping the template with the refined homography, projecting it onto the original template, and calculating the IoU.
@@
Projection Error: The projection error was quantified as the average distance, in meters, between the projected points using the estimated homography and the corresponding GT. To achieve that, we uniformly sampled 2,500 pixels from the visible field area of the camera image and projected them onto the field to compute the distance. The standard dimensions of a soccer field are 105$\times$68 meters.
@@
Reprojection Error: The reprojection error was calculated by averaging the distance between the reprojected points in the video frame, utilizing both the estimated and the GT homography.
IV-C Implementation Details
@@
Due to the absence of publicly available results on the multiple-view SN23 [8] distribution, we trained two models from scratch: Multi-view (MV) and Single-view (SV). The latter is trained on a data subset composed almost entirely of non-replay frames, ensuring a high percentage of central camera shots. We train separate networks for the keypoints and line extremities detection tasks on the SN23-train dataset [8]. For the MV model, we train for $200$ epochs, using an initial learning rate of $1e^{-2}$ and a batch size of $2$. For the SV model, we train for $200$ epochs, using an initial learning rate of $1e^{-5}$ and a batch size of $1$. We utilize the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. ${l}_{2}$-norm loss is used for heatmap regression in both neural networks. Data augmentation such as random horizontal flip, color jitter, and Gaussian noise are applied to enhance model robustness and generalization. Furthermore, we fine-tune both SV networks, keypoint and line extremity detection, on the WC14 [23] and TSWC [7] datasets. GT homographies are transformed into our proposed keypoint sets and line extremities by projecting their respective world coordinates to the field’s ground plane. Note that non-planar points and lines are excluded from the transformation process, and their corresponding output layers are masked in the loss function. In fine-tuning, input images are resized to match the SoccerNet dataset image size. The experiments are conducted on a single NVIDIA GeForce RTX 2080 Ti GPU with $12$ GB of memory, and the implementation is carried out in the PyTorch framework. Calibration and further optimization tasks are conducted on a Intel Xeon Silver 4214 Processor and the implementations are carried out in the opencv-python and scipy frameworks, respectively.
IV-D Results and comparisons
@@
We now present the results of an extensive evaluation, divided into camera calibration and homography estimation. The camera calibration evaluation assesses the accuracy of individual camera parameters using the $\text{JaC}_{\gamma}$ metric, while the homography estimation is evaluated using the IoU one, projection error, and reprojection error. As previously noted, methods with the subscript MV and SV correspond to multi-view and single-view models, respectively, and the PnL designation indicates the use of point and line calibration refinement module.
IV-D1 Camera Calibration
@@
In team sports such as soccer, the action takes place on a nearly planar field. Consequently, most methods utilize homography estimation to map all elements positioned on this plane but cannot project non-planar points such as points belonging to goal posts or crossbars. Conversely, in [46, 32, 14] make use of a 3D model of the soccer field to extract camera pose and intrinsic parameters directly. In homography-based approaches, parameter retrieval is accomplished through homography decomposition (HDecomp). We conduct a quantitative comparison of our proposed method to state-of-the-art approaches [5, 26, 46, 32] on the SN22-test-center dataset, comprising only images where the main camera center is visible (1,454 images). Furthermore, utilizing the SoccerNet annotation format for the WC14-test dataset provided by Theiner et al. [46], we conduct a comparison of our proposed method’s performance in camera parameter estimation on the WorldCup 2014 dataset distribution.
We report the statistics from [46] for the results of state-of-the-art approaches [5, 26, 46], along with results provided by [32]. As it is shown in Tables I-II for the SN22-test and WC14 datasets, respectively, our SV method outperforms state-of-the-art approaches across several metrics in both datasets. Additionally, the inclusion of the PnL module demonstrates its effectiveness by increasing the FS metric by $4.3\%$ on the SN22-test-center dataset and a $7.6\%$ increase on the WC14-test one. Minor variations in CR, compared to previous approaches, stem from our method’s requirement for a minimum number of visible keypoints for calibration and differences in the maximum allowable reprojection error to consider the parameter estimation valid. Moreover, variations in CR between our SV models are also due to the differing maximum allowable reprojection errors. Reprojection error is computed based on keypoint errors, meaning that while line-based refinement increases keypoint reprojection error, it results in improved overall calibration. Qualitative results demonstrating the effect of the PnL refinement module are presented in Fig. 4.
$\text{JaC}_{\gamma}$ [%]
@@
SN23-test [14] 76.6 - - 73.6 56.3
[…]
TABLE I: Evaluating camera calibration on SoccerNet [12] distributions. The table reports our results for the Single-view model on the SN22-test-center dataset as well as for the Multi-view model on the full SN23-test dataset, including in both cases comparisons with competing approaches. For computing the $\text{JaC}_{\gamma}$ metric, we consider $\gamma=\{5,10,20\}$.
$\text{JaC}_{\gamma}$ [%]
@@
[32] 59.4 82.3 90.9 100.0 59.4
@@
TABLE II: Quantitative comparison of our Single-view model on camera calibration conducted on the WC14-test dataset. For computing the $\text{JaC}_{\gamma}$ metric, we consider $\gamma=\{5,10,20\}$.
We also evaluate our method on the entire SN23-test dataset, reporting statistics from [14], which represents the only comparable approach capable of extending calibration assessments to full multi-view scenarios. Although the calibration approach proposed by Mavrogiannis et al. [32] extends beyond the main camera, it is limited to calibrating the main, offside, and behind-the-goalpost cameras. Moreover, their method requires training a separate model for each camera location, whereas our approach generalizes keypoint and line extremity detection across the entire SoccerNet distribution using a single model. As shown in Table I for the SN23-test datasets, our MV model outperforms the existing approaches across all the presented metrics. Additionally, our PnL refinement module showcases a $2.8\%$ increase in the FS metric. The cost function’s weighting parameter is empirically set to $\alpha=0.5$ to maximize the FS metric. Qualitative results showcasing calibration results of our base pipeline on different camera angles are presented in Fig. 5. The significant decrease in performance compared to the SN22-test-center dataset is attributed to the challenges of calibrating images with camera poses that deviate substantially from the main camera, such as close-up shots, where few or no landmarks are visible, or fisheye shots from inside the goals, where extreme lens distortion makes calibration unable. Examples of these challenging camera shots are shown in Fig. 6.
In terms of inference speed, the plain MV model achieves 7 Hz by using the proposed image size and the software and hardware configurations described above. When incorporating the PnL refinement module, the inference speed decreases to an average of 4 Hz during evaluation on the SN23-test dataset. However, the speed for specific frames varies significantly, depending on the number of detected lines and the initialization quality.
Figure 4: Qualitative results of the PnL refinement module. Top: Projection of soccer field lines with initial calibration estimate $\{\mathbf{K},\mathbf{R},\mathbf{t}\}$. Bottom: Projection of soccer field lines with refined calibration $\{\mathbf{K},\mathbf{R^{\prime}},\mathbf{t^{\prime}}\}$ through PnL refinement module.
IV-D2 Homography Estimation
The proposed method is compared with respect to state-of-the-art approaches [5, 26, 9, 52, 53, 33, 40, 7, 34, 46, 29] using the WC14-test dataset. Additionally, our method is also compared to state-of-the-art approaches [5, 33, 7, 29, 34] using the TSWC-test dataset. For computing IoU-based metrics, projection error, and reprojection error, we adopt the approach outlined in [7]. The dimensions of the soccer field template are set at $115\times 74$ yards for a fair comparison with previous approaches.
In the evaluation on the WC14-test dataset, we report performance metrics based on the findings from the respective works, as shown in Table III. For a fair comparison with [9], we use the results from their ours-w/o-players approach, as reported in their paper. Our fine-tuned model produces comparable results to previous approaches across all IoU-based, projection error, and reprojection error metrics. However, with the addition of the PnL refinement module, our method achieves state-of-the-art solutions across all metrics, further demonstrating the effectiveness of the proposed refinement module also in the homography estimation task.
For the evaluation on the TSWC-test dataset, we report the results for [33, 7, 29, 34] in Table III, observing how our fine-tuned model with the PnL refinement module outperforms those methods across all metrics.
Dataset Approach $\text{IoU}_{\text{part}}$uparrow (%) $\text{IoU}_{\text{whole}}$uparrow (%) Proj. $\downarrow$ (m) Reproj. $\downarrow$
Mean Median Mean Median Mean Median Mean Median
@@ IV-D2 Homography Estimation
TABLE III: Evaluating the homography estimation on WC14-test and TSWC-test. $\ast$ denotes the methods trained on SoccerNet distribution, $\dagger$ denotes the methods fine-tuned on the WC14 dataset and $\ddagger$ denotes the methods fine-tuned on the TSWC one.
IV-D3 Ablation Study on Keypoint Sets Contribution
The contribution of each keypoint set, namely $\mathcal{K}p_{e}$, $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$, and $\mathcal{K}p_{3}$, is analyzed in Table IV. Overall, the integration of each keypoint set results in improvements in the CR and FS metrics, with each set contributing to different geometric elements of the field. The inclusion of $\mathcal{K}p_{e}$ increases CR by providing visible landmarks along the main field’s straight lines, particularly when line-line intersections are scarce in the image. However, despite these improvements in CR and FS, the Acc metrics decrease, as the method is able to calibrate more lines but with some lower-quality calibrations. A similar pattern is observed with the integration of the $\mathcal{K}p_{1}$ set, which boosts CR and FS by adding more visible landmarks to the field’s straight lines and enabling the inclusion of field circles, though calibration quality in these areas remains suboptimal. A significant increase in CR, Acc, and consequently FS, is achieved through the integration of the $\mathcal{K}p_{2}$ set, particularly in frames where the field’s center circle is partially visible, but midfield line intersections are absent from the image. Finally, the integration of the $\mathcal{K}p_{3}$ set leads to further improvements across all metrics, enhancing the robustness of the method and setting a new state-of-the-art in sports field registration benchmarks.
$\text{JaC}_{\gamma}$ [%]
$\mathcal{K}p_{e}$ $\mathcal{K}p_{1}$ $\mathcal{K}p_{2}$ $\mathcal{K}p_{3}$ 5 10 20 CR FS
74.2 89.4 93.7 86.1 63.9
74.0 89.0 93.5 88.8 65.7
74.0 88.3 91.9 94.4 69.9
75.1 89.2 92.6 97.8 73.5
75.8 89.7 91.9 98.1 74.4
@@ IV-D3 Ablation Study on Keypoint Sets Contribution
TABLE IV: Ablation study of our keypoint sets. The table shows the effect of every keypoint set on the SN22-test-center dataset.
Figure 6: Challenging camera shots where calibration is unable. Left: Fisheye shots from inside the goals with extreme lens distortion. Right: Close-up shots with few or none visible landmarks.
IV-D4 Ablation Study on Refinement Contribution
The contribution of each field geometrical object for initial estimate refinement—namely points, lines, and the full PnL module—is analyzed in Table V. Similar behaviors are observed across the SoccerNet distributions: after obtaining the initial calibration estimate using the proposed keypoint sets, point-based refinement does not significantly alter the results. While point refinement slightly reduces reprojection error, leading to a minor increase in the CR metric, it simultaneously lowers accuracy metrics. In contrast, line-based refinement not only reduces reprojection error but also improves accuracy metrics, as reprojection error is calculated using points, and the maximum allowable reprojection error in the method affects these outcomes. Thus, line refinement generally produces more accurate calibration results. Finally, we confirm the previously noted effectiveness of the PnL module, which balances the contributions of both geometrical objects to deliver state-of-the-art results in the SoccerNet benchmarks.
For the WC14 dataset, a different behavior is observed: point-based refinement has no impact, and line-based refinement results in worse calibration across all metrics. Nevertheless, the proposed PnL refinement still achieves state-of-the-art results.
Dataset $\text{JaC}_{\gamma}$ [%]
P L 5 10 20 CR FS
SN23-test 72.2 84.9 88.4 80.4 58.1
71.6 84.2 87.7 81.3 58.2
74.3 86.4 90.8 74.9 55.7
76.7 87.2 90.1 79.5 60.9
SN22-test -center 75.8 89.7 91.9 98.1 74.4
75.8 89.8 93.7 98.1 74.4
76.3 89.4 94.3 96.5 73.7
80.6 89.9 92.4 97.7 78.7
WC14-test 77.6 89.8 93.7 100.0 77.6
77.6 89.8 93.7 100.0 77.6
76.5 90.8 94.5 99.4 76.1
85.2 94.0 96.1 100.0 85.2
TABLE V: Ablation study of our keypoint sets. The table shows the effect of every keypoint set on the SN22-test-center dataset.
V Conclusion
@@
In this paper, we introduce a novel framework for 3D sports field registration. Our proposed pipeline adopts a minimalist approach by solely utilizing the geometric properties of the soccer field. We demonstrate superior performance in 3D camera calibration on SoccerNet and WorldCup 2014 datasets compared to state-of-the-art methods, while also achieving comparable results in homography estimation on WorldCup 2014 and TS-WorldCup datasets. Additionally, we introduce a novel calibration refinement module that leverages field landmarks, such as keypoints and lines, which, when integrated with our calibration pipeline, further enhance performance in 3D camera calibration benchmarks and achieve state-of-the-art results in homography estimation. We also extend our pipeline to multi-view camera calibration, establishing the state-of-the-art benchmark for multi-view broadcast-based camera calibration in soccer. Our method exhibits promising results, highlighting the effectiveness of utilizing a robust field template without the need for further refinements. Furthermore, the refinement module showcases its effectiveness showing superior performance and improving calibration results when visible field landmarks are scarce. As long as the video frame distortions are not too harsh (i.e., fisheye camera shots) and a minimum of four keypoints are visible, the proposed sports field registration approach is shown to be effective. In future work, we plan to enhance our approach by incorporating temporal consistency between subsequent video frames, aligning better with the nature of sports video broadcasts.
@@
Acknowledgment. This work has been supported by the project GRAVATAR PID2023-151184OB-I00 funded by MCIU/AEI/10.13039/501100011033 and by ERDF, UE and by the Government of Catalonia under Joan Oró FI 2024 grant.